# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

### v1

@@ 3.4 Koopman-guided spectrum evolvement
Motivated by Koopman operator theory [29, 40, 5], we introduce a layer to capture the temporal evolution of time-frequency patterns in wavelet space. We aim to learn a Koopman operator with $K$ dimensions in the transformed space. To obtain the operator $\mathbf{K}\in\mathbb{C}^{K\times K}$, we first initialise a learnable complex-valued matrix $\mathbf{S}\in\mathbb{C}^{K\times K}$. At each forward pass, we apply QR decomposition to $\mathbf{S}$, and retain only the resulting unitary matrix $\mathbf{U}\in\mathbb{C}^{K\times K}$, i.e. $\mathbf{U}\leftarrow\texttt{QR}\left(\mathbf{S}\right)$ s.t. $\mathbf{U}^{\dagger}\mathbf{U}=\mathbf{I}$, where $\mathbf{U}^{\dagger}$ is the conjugate transpose of $\mathbf{U}$. Multiplying with $\mathbf{U}$ therefore prevents data amplification or distortion.
We then initialise a learnable vector $\mathbf{p}\in\mathbb{R}^{K}$, where the $k$-th element, $p_{k}$, controls the temporal evolution (phase transformation angle) of the $k$-th transformation. All elements in the vector are then mapped into complex numbers, obtaining $\mathbf{v}\in\mathbb{C}^{K}$, where $v_{k}=e^{ip_{k}}$. We use $\mathbf{v}$ to form a diagonal matrix $\mathbf{D}\in\mathbb{C}^{K\times K}=\texttt{diag}(\mathbf{v})$. The Koopman operator, $\mathbf{K}$, is then given by $\mathbf{K}=\mathbf{U}\mathbf{D}\mathbf{U}^{\dagger}$. We use $\mathbf{K}$ to model the temporal evolution of the MVCA embedding, $\mathbf{O}$, after converting it to a complex form, $\mathbf{O}_{c}\in\mathbb{C}^{K\times L\times d}$ (imaginary part is set to $i0$):
$\mathbf{O}_{l}=\mathbf{K}\times\mathbf{O}_{c}\,\,,$ (4)
where $\mathbf{O}_{l}\in\mathbb{C}^{K\times L\times d}$ is the evolved embedding after multiplication with $\mathbf{K}$.
A conventional Koopman framework models temporal evolution recursively at each time step [28, 5]. We instead choose to apply Koopman transformation in the frequency domain with one forward pass, which is equivalent to learning a direct transformation from the input to the output. This global projection reduces the accumulation of sequential errors while maintaining robustness in training.
@@ 3.6 Convolutional decoder
Finally, a $3$-layer convolutional decoder transforms the learned representation. It comprises $3$ $1$-dimensional convolutional layers with GELU activations between every $2$ layers. Each layer uses kernel sizes $[5,3,3]$ and paddings $[2,1,1]$ respectively, followed by an adaptive average pooling layer at the end. The dimension of each layer is $[H\times 4,\,H\times 2,\,H]$, producing the final sequence representation $\mathbf{Z}_{\text{out}}\in\mathbb{R}^{H\times H}$, in accordance with the time steps of the target forecasting horizon. The result is then linearly projected using a weight vector $\mathbf{W}_{z}\in\mathbb{R}^{H}$ to generate the final output, as in $\hat{\mathbf{y}}\in\mathbb{R}^{H}=\mathbf{Z}_{\text{out}}\mathbf{W}_{z}$, in accordance with the dimensionality of the target variable.
@@ 4 Results
We assess forecasting accuracy using an expanded collection of data sets and tasks, to overcome potential biases present in the current literature. We first compare Sonnet against other competitive baseline models. We then investigate the role of attention mechanisms in time series forecasting models by removing or replacing naïve transformers with more advanced variants, including the proposed MVCA module. We also provide an ablation study of the key components of Sonnet.
4.1 Experiment settings
We conduct experiments over $12$ real-world data sets. This includes $2$ established benchmarks from prior papers [61, 58], specifically the ETTh1 and ETTh2 data sets, which contain hourly electricity transformer temperature forecasting. Oil temperature is our target variable, with the remaining indicators considered as exogenous variables. We also use $2$ data sets from the Darts library [16], predicting hourly energy prices (ENER) and electricity (low-voltage) consumption (ELEC). In addition, we form weather data sets extracted from the WeatherBench repository [38] for $5$ cities from diverse geographical locations—London (WEA-LD), New York (WEA-NY), Hong Kong (WEA-HK), Cape Town (WEA-CT), and Singapore (WEA-SG)—to support a more inclusive analysis. For each city, we sample data from its nearest grid point and obtain $5$ climate indicators. We provide spatial context by including data from its eight surrounding grid points ($3\times 3$ grid) as additional exogenous variables. Following prior work on global climate forecasting [49], we resample the data to a temporal resolution of $6$ hours. The forecasting target is the $850$ hPa (T850) temperature, a key indicator for climate modelling [41, 14]. Finally, we include influenza-like illness (ILI) rate forecasting tasks (as in [43]) in $3$ locations, England (ILI-ENG), and in U.S. Health & Human Services (HHS) Regions 2 (ILI-US2) and 9 (ILI-US9). In the ILI tasks, frequency time series of web searches are included as exogenous predictors. More information about the data sets is provided in Appendix A.
For the ETT tasks, we set the forecasting horizon ($H$) to $\{96,192,336,720\}$ time steps, and use a single test set of consecutive unseen instances adopting the evaluation settings in prior work [34, 25]. For ENER we set $H=\{24,48,72,168\}$ hours ahead and use $1$ year (2018) for testing. For ELEC, we use the last $2$ years (2020, ’21) as $2$ distinct test seasons, and set $H$ to $\{12,24,36\}$ hours. For the WEA tasks, we form $3$ test sets (years 2016, ’17, ’18), and set $H$ to $\{4,12,28,120\}$ time steps corresponding to $\{1,3,7,30\}$ days. Following the same setup as in [43], for the ILI forecasting task, we test models on $4$ consecutive influenza seasons (2015/16 to 2018/19), each time training a new model on data from previous seasons. We set $H=\{7,14,21,28\}$ days. Further details are provided in Appendix D.
We compare Sonnet to $8$ competitive forecasting models that, to the best of our knowledge, form the current SOTA methods: DLinear [58], Crossformer [60], iTransformer [25], PatchTST [34], TimeXer [52], Samformer [18], ModernTCN [27], and DeformTime [43]. We also include a naïve baseline which can either be a seasonal or a standard persistence model, depending on the presence of strong seasonality. For all tasks (except ETT), we conduct hyperparameter tuning for all models. For the ETT tasks, we adopt settings from the official repositories of methods (expect for Crossformer that did not provide this). Further details about the baseline models can be found in Appendix C.
Table 1: Forecasting accuracy results across all tasks, methods, and forecasting horizons ($H$). For the ELEC/ILI/WEA tasks, we report the average performance across $2$, $4$, and $3$ test sets, respectively (detailed breakdowns in Appendix E.1). $\epsilon\%$ denotes sMAPE. Best results are bolded and second best are underlined. Grey background denotes instances where a model does not outperform persistence.
@@ 4.1 Experiment settings
4.2 Forecasting accuracy of Sonnet
Prediction accuracy for all tasks is enumerated in Table 1, using Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage Error (sMAPE or $\epsilon\%$) as the main evaluation metrics. For WEA tasks, sMAPE is modified to avoid small values due to Kelvin temperature units (see Appendix D.6). For the ETT and ENER tasks, results are based on a single forecasting season (i.e., one train / test split per forecasting horizon), while for other tasks we average over multiple seasons.
Sonnet exhibits the overall best performance across the explored MTS forecasting tasks. Comparing Sonnet’s performance to the best-performing baseline model for each task and forecasting horizon yields an MAE reduction of $1.1\%$ on average. Ranking-wise, Sonnet is the best-performing model on $34$, and the second-best on $9$ out of $47$ forecasting tasks. Sonnet shows relatively inferior performance on the ETT tasks. However, the ETT data sets have only $6$ exogenous variables and cover a limited training time span ($2$ years). Consequently, they offer a limited spectrum to gain from exploring variable dependencies, as well as insufficient historical context, both of which are elements that Sonnet leverages from. Aside from the ETT tasks, Sonnet reduces MAE by $3.3\%$ on average. Therefore, Sonnet is more effective with more covariates and longer time spans for training. Compared to a specific forecasting model, Sonnet consistently improves accuracy, reducing MAE from $2.2\%$ for the best-performing baseline (DeformTime) to $31.1\%$ for the worst-performing one (DLinear). This further highlights the model’s forecasting capacity amongst a wide range of applications.
Focusing on the more challenging forecasting tasks (ILI and WEA), Sonnet outperforms baselines $25$ or $26$ (out of $32$ tasks in total) based on sMAPE or MAE, respectively. The MAE reduction is $3.5\%$ on average for ILI tasks and $2\%$ for WEA tasks. Within these tasks, we also observe consistent comparative performance patterns in the baseline models. Models that do not capture inter-variable dependencies (DLinear and PatchTST) generally offer lower predictability (Sonnet reduces their MAE by $28.8\%$ and $29.8\%$), and in most ILI tasks, they cannot surpass the performance of a persistence model. In contrast, baseline models that capture inter-variable dependencies tend to perform better. Among these, those that embed the covariates along the temporal dimension—namely Samformer, TimeXer, and iTransformer—exhibit inferior performance (Sonnet reduces their MAE by $23.9\%$, $22.1\%$, and $22.9\%$), whereas models that preserve temporal order—namely DeformTime, ModernTCN, and Crossformer—achieve stronger results (Sonnet reduces their MAE by $3.4\%$, $11.1\%$, and $8\%$). Hence, in tasks with more informative covariates, preserving temporal structure while modelling inter-variable dependencies is a desirable property.
Table 2: Performance (average MAE across the $4$ test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. ‘$\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet [43], FEDformer [62], and DeformTime [43], respectively. Best results are bolded and second best are underlined.
@@ 4.3 Effectiveness of different attention modules in MTS forecasting
We evaluate the effectiveness of the proposed attention-driven module, MVCA, by integrating it into existing forecasting models (we refer to them as base models) that originally deploy naïve transformer attention. Experiments are conducted on the ILI tasks as these are the most representative of the MTS class we are exploring: they not only contain many exogenous predictors, but also frame hard practical epidemiological modelling problems when $H=21$ or $28$ days ahead. We use Samformer [18], iTransformer [25], and PatchTST [34] as base models. For each one of them, we evaluate $5$ attention configurations: removing the attention module altogether ($\neg$ Attn), attention with Fourier transformer proposed by FNet [21], the Frequency Enhanced Decomposed (FED) attention proposed in FEDformer [62], the variable deformable attention (VDAB) proposed by DeformTime [43], and ours (MVCA). Further details are provided in Appendices C.2 and D.8.
MAEs averaged across the $4$ test seasons for all forecasting horizons are enumerated in Table 2 (sMAPEs are reported in Appendix E.2, Table S9). Notably, for the base models, removing the attention module does not lead to significant performance degradation, with some cases even showing improved results, e.g. improved accuracy is reported for some horizons for the ILI-ENG task. This indicates that the attention modules in the base models do not always capture useful information.

### v2

@@ 3.4 Koopman-Guided Spectrum Evolvement
Motivated by Koopman operator theory (mezic2005spectral; rowley2009spectral; avila2020datadriven), which offers a framework for modelling nonlinear dynamics, we introduce a layer to capture the temporal evolution of time-frequency patterns in wavelet space. We aim to learn a Koopman operator with $K$ dimensions in the transformed space. To obtain the operator $\mathbf{K}\in\mathbb{C}^{K\times K}$, we first initialise a learnable complex-valued matrix $\mathbf{S}\in\mathbb{C}^{K\times K}$. At each forward pass, we apply QR decomposition to $\mathbf{S}$, and retain only the resulting unitary matrix $\mathbf{U}\in\mathbb{C}^{K\times K}$, i.e. $\mathbf{U}\leftarrow\mathrm{QR}\left(\mathbf{S}\right)$ s.t. $\mathbf{U}^{\dagger}\mathbf{U}=\mathbf{I}$, where $\mathbf{U}^{\dagger}$ is the conjugate transpose of $\mathbf{U}$. Multiplying with $\mathbf{U}$ therefore prevents data amplification or distortion.
We then initialise a learnable vector $\mathbf{p}\in\mathbb{R}^{K}$, where the $k$-th element, $p_{k}$, controls the temporal evolution (phase transformation angle) of the $k$-th transformation. All elements in the vector are then mapped into complex numbers, obtaining $\mathbf{v}\in\mathbb{C}^{K}$, where $v_{k}=e^{ip_{k}}$. We use $\mathbf{v}$ to form a diagonal matrix $\mathbf{D}\in\mathbb{C}^{K\times K}=\mathrm{diag}(\mathbf{v})$. The Koopman operator, $\mathbf{K}$, is then given by $\mathbf{K}=\mathbf{U}\mathbf{D}\mathbf{U}^{\dagger}$. We use $\mathbf{K}$ to model the temporal evolution of the MVCA embedding, $\mathbf{O}$, after converting it to a complex form, $\mathbf{O}_{c}\in\mathbb{C}^{K\times L\times d}$ (the imaginary part is set to $i0$ to perform complex-valued transformations without altering the original embedding), given by $\mathbf{O}_{l}=\mathbf{K}\times\mathbf{O}_{c}$, where $\mathbf{O}_{l}\in\mathbb{C}^{K\times L\times d}$ is the evolved embedding after multiplication with $\mathbf{K}$. A conventional Koopman framework models temporal evolution recursively at each time step (Lusch2018deep; avila2020datadriven). We instead choose to apply Koopman transformation in the frequency domain with one forward pass, which is equivalent to learning a direct transformation from the input to the output. This global projection reduces the accumulation of sequential errors while maintaining robustness in training.
@@ 3.6 Convolutional Decoder
Finally, a $3$-layer convolutional decoder transforms the learned representation. It comprises 33 11-dimensional convolutional layers with GELU activations between every $2$ layers. Each layer uses kernel sizes $[5,3,3]$ and paddings $[2,1,1]$ respectively, followed by an adaptive average pooling layer at the end. The dimension of each layer is $H\times 4,\,H\times 2,\,H]$, producing the final sequence representation $\mathbf{Z}_{\text{out}}\in\mathbb{R}^{H\times H}$, in accordance with the time steps of the target forecasting horizon. The result is then linearly projected using a weight vector $\mathbf{W}_{z}\in\mathbb{R}^{H}$ to generate the final output, as in $\hat{\mathbf{y}}\in\mathbb{R}^{H}=\mathbf{Z}_{\text{out}}\mathbf{W}_{z}$, in accordance with the dimensionality of the target variable.
@@ 4 Results
We assess forecasting accuracy using an expanded collection of data sets and tasks, to overcome potential biases present in the current literature. We first compare Sonnet against other competitive baseline models. We then investigate the role of attention mechanisms in time series forecasting models by removing or replacing naïve transformers with more advanced variants, including the proposed MVCA module. We also provide an ablation study of the key components of Sonnet and seed control in Appendix E.
@@
Table 1: Forecasting accuracy results across all tasks, methods, and forecasting horizons ($H$). For the ELEC/ILI/WEA tasks, we report the average performance across 22, 44, and $3$ test sets, respectively (detailed breakdowns in Appendix E.1). $\epsilon\%$ denotes sMAPE. Best results are bolded, second best underlined. Grey background denotes the model does not outperform persistence.
4.1 Experiment Settings
We conduct experiments over $12$ real-world data sets. This includes $2$ established benchmarks from prior papers (zhou2021informer; zeng2023are), specifically the ETTh1 and ETTh2 data sets, which contain hourly electricity transformer temperature forecasting. Oil temperature is our target variable, with the remaining indicators considered as exogenous variables, following wang2024timexer. We also use $2$ data sets from the Darts library (herzen2022darts), predicting hourly energy prices (ENER) and electricity (low-voltage) consumption (ELEC). In addition, we form weather data sets extracted from the WeatherBench repository (rasp2020weatherbench) for $5$ cities from diverse geographical locations, namely London (WEA-LD), New York (WEA-NY), Hong Kong (WEA-HK), Cape Town (WEA-CT), and Singapore (WEA-SG), to support a more inclusive analysis. For each city, we sample data from its nearest grid point and obtain $5$ climate indicators. We provide spatial context by including data from its eight surrounding grid points ($3\times 3$ grid) as additional exogenous variables. Following prior work on global climate forecasting (verma2024climode), we resample the data to a temporal resolution of $6$ hours. The forecasting target is the $850$ hPa (T850) temperature, a key indicator for climate modelling (scherrer2004analysis; hamill2007ensemble). Finally, we include influenza-like illness (ILI) rate forecasting tasks (as in (shu2025deformtime)) in $3$ locations, England (ILI-ENG), and in U.S. Health & Human Services (HHS) Regions 2 (ILI-US2) and 9 (ILI-US9). In the ILI tasks, frequency time series of web searches are included as exogenous predictors. More information about the data sets is provided in Appendix A.
For the ETT tasks, we set the forecasting horizon ($H$) to $\{96,192,336,720\}$ time steps, and use a single test set of consecutive unseen instances adopting the evaluation settings in prior work (nie2023time; liu2024itransformer). For ENER we set $H=\{24,48,72,168\}$ hours ahead and use $1$ year (2018) for testing. For ELEC, we use the last $2$ years (2020, ’21) as $2$ distinct test seasons, and set $H$ to $\{12,24,36\}$ hours. For the WEA tasks, we form $3$ test sets (years 2016, ’17, ’18), and set $H$ to $\{4,12,28,120\}$ time steps corresponding to $\{1,3,7,30\}$ days. Following the same setup as in (shu2025deformtime), for the ILI forecasting task, we test models on $4$ consecutive influenza seasons (2015/16 to 2018/19), each time training a new model on data from previous seasons. We set $H=\{7,14,21,28\}$ days (more details in Appendix D).
We compare Sonnet to $8$ competitive forecasting models that, to the best of our knowledge, form the current SOTA methods: DLinear (zeng2023are), Crossformer (zhang2023crossformer), iTransformer (liu2024itransformer), PatchTST (nie2023time), TimeXer (wang2024timexer), Samformer (ilbert2024samformer), ModernTCN (luo2024moderntcn), and DeformTime (shu2025deformtime), with TimeXer and DeformTime being designed for multivariable forecasting. We also include a naïve baseline which can either be a seasonal or a standard persistence model, depending on the presence of strong seasonality. For all tasks (except ETT), we conduct hyperparameter tuning for all models. For the ETT tasks, we adopt settings from the official repositories of methods (except for Crossformer, which did not provide this). Appendix C has further details about the baseline models.
@@ 4.1 Experiment Settings
Table 2: Performance (average MAE across 44 test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. $\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet (shu2025deformtime), FEDformer (zhou2022fedformer), and DeformTime (shu2025deformtime), respectively. Best results are bolded, second best underlined.
4.2 Forecasting Accuracy of Sonnet
Prediction accuracy for all tasks is enumerated in Table 1, using Mean Absolute Error (MAE) and symmetric Mean Absolute Percentage Error (sMAPE or $\epsilon\%$) as the main evaluation metrics. For WEA tasks, sMAPE is modified to avoid small values due to Kelvin temperature units (see Appendix D.6). For ETT and ENER tasks, results are based on a single forecasting season (i.e., one train / test split per forecasting horizon), while we average over multiple seasons for other tasks.
Sonnet exhibits the overall best performance across the explored MTS forecasting tasks. Comparing Sonnet’s performance to the best-performing baseline model for each task and forecasting horizon yields an MAE reduction of $1.1\%$ on average. Ranking-wise, Sonnet is the best-performing model on $34$, and the second-best on $9$ out of $47$ forecasting tasks. Sonnet shows relatively inferior performance on the ETT tasks. However, the ETT data sets have only $6$ exogenous variables and cover a limited training time span ($2$ years). Consequently, they offer a limited spectrum to gain from exploring variable dependencies, as well as insufficient historical context, both of which are elements that Sonnet leverages from. Aside from the ETT tasks, Sonnet reduces MAE by $3.3\%$ on average. Therefore, Sonnet is more effective with more covariates and longer time spans for training. Compared to a specific forecasting model, Sonnet consistently improves accuracy, reducing MAE from $2.2\%$ for the best-performing baseline (DeformTime) to $31.1\%$ for the worst-performing one (DLinear). We note that the MAE performance gain over DeformTime is statistically significant based on a paired t-test across all the $47$ forecasting tasks (p=5e4p=5e-4). This further highlights the model’s forecasting capacity amongst a wide range of applications.
Focusing on the more challenging forecasting tasks (ILI and WEA), Sonnet outperforms baselines $25$ or $26$ (out of $32$ tasks in total) based on sMAPE or MAE, respectively. The MAE reduction is $3.5\%$ on average for ILI tasks and $2\%$ for WEA tasks. Within these tasks, we also observe consistent comparative performance patterns in the baseline models. Models that do not capture inter-variable dependencies (DLinear and PatchTST) generally offer lower predictability (Sonnet reduces their MAE by $28.8\%$ and $29.8\%$), and in most ILI tasks, they cannot surpass the performance of a persistence model. In contrast, baseline models that capture inter-variable dependencies tend to perform better. Among these, those that embed the covariates along the temporal dimension, namely Samformer, TimeXer, and iTransformer, exhibit inferior performance (Sonnet reduces their MAE by $3.9\%, 22$$2.1\%$, and $22.9\%$), whereas models that preserve temporal order, namely DeformTime, ModernTCN, and Crossformer, achieve stronger results (Sonnet reduces their MAE by $3.4\%$, $11.1\%$, and $8\%$). Hence, in tasks with more informative covariates, preserving temporal structure while modelling inter-variable dependencies is a desirable property. We provide ablation study results in Appendix E.3 that quantify the contribution of different modules to the overall forecasting accuracy of Sonnet. Additional performance evaluation results over the entire output sequence are provided in Appendix E.8.
@@ 4.3 Effectiveness of Different Attention Modules in MTS Forecasting
We evaluate the effectiveness of the proposed attention-driven module, MVCA, by integrating it into existing forecasting models (we refer to them as base models) that originally deploy naïve transformer attention. Experiments are conducted on the ILI tasks as these are the most representative of the MTS class we are exploring: they not only contain many exogenous predictors, but also frame hard practical epidemiological modelling problems when $H=21$ or $28$ days ahead. We use Samformer (ilbert2024samformer), iTransformer (liu2024itransformer), and PatchTST (nie2023time) as base models. For each one of them, we evaluate $5$ attention configurations: removing the attention module altogether ($\neg$ Attn), attention with Fourier transformer proposed by FNet (lee-thorp2022fnet), the Frequency Enhanced Decomposed (FED) attention proposed in FEDformer (zhou2022fedformer), the variable deformable attention (VDAB) proposed by DeformTime (shu2025deformtime), and ours (MVCA). Further details are provided in Appendices C.3 and D.8.
Table 2 enumerates MAEs averaged across the $4$ test seasons for all forecasting horizons (sMAPEs in Appendix E.2, Table S9). Notably, for the base models, removing the attention module does not lead to significant performance degradation, with some cases even showing improved results, e.g. improved accuracy is reported for some horizons for the ILI-ENG task. This indicates that the attention modules in the base models do not always capture useful information.
Sonnet VDAB FED FNet Naïve
ILI-ENG
7 1.479 21.8 1.433 20.3 5.991 77.7 1.591 27.7 1.717 25.3
14 1.923 25.8 1.973 29.4 6.031 78.1 2.084 32.8 2.068 31.4
21 2.510 36.5 2.774 35.2 6.045 78.2 2.765 39.3 2.781 37.2
28 2.748 37.0 3.060 44.3 6.051 78.4 3.240 44.5 3.078 50.0
ILI-US2
7 0.381 14.9 0.429 16.6 1.168 45.0 0.415 18.5 0.436 17.5
14 0.449 18.4 0.500 19.1 1.179 45.5 0.558 23.2 0.532 20.4
21 0.533 20.6 0.575 20.5 1.195 46.2 0.659 25.3 0.613 23.5
28 0.579 21.1 0.643 24.3 1.193 46.1 0.705 29.1 0.651 26.1
ILI-US9
7 0.267 13.0 0.277 13.4 0.856 41.0 0.286 14.4 0.301 13.9
14 0.281 13.1 0.324 15.8 0.856 41.0 0.351 16.1 0.340 14.9
21 0.318 14.1 0.319 14.5 0.857 41.0 0.379 17.8 0.404 19.2
28 0.367 16.5 0.375 16.8 0.857 41.0 0.379 18.2 0.411 19.1
WEA-CT
4 1.624 9.6 1.714 10.1 4.359 24.6 1.629 9.6 1.726 10.2
12 3.543 20.4 3.563 20.5 4.359 24.7 3.588 20.6 3.558 20.5
28 3.728 21.3 3.760 21.5 4.359 24.6 3.916 22.3 3.758 21.5
120 3.737 21.4 3.734 21.3 4.356 24.6 3.815 21.7 3.763 21.5
WEA-HK
4 0.639 4.1 0.645 4.1 2.801 17.7 0.639 4.1 0.676 4.3
12 1.235 7.7 1.279 8.0 2.801 17.7 1.325 8.3 1.281 8.1
28 1.413 8.8 1.463 9.1 2.808 17.8 1.494 9.3 1.475 9.2
120 1.547 9.6 1.547 9.6 2.810 17.8 1.639 10.1 1.621 10.0
WEA-LD
4 1.723 15.2 1.814 15.8 4.668 34.9 1.740 15.3 1.811 16.0
12 2.959 23.9 2.986 24.1 4.661 34.9 3.008 24.1 2.973 24.0
28 3.216 25.5 3.271 25.8 4.659 34.9 3.378 26.5 3.268 25.8
120 3.246 25.8 3.319 26.2 4.668 34.9 3.667 28.5 3.407 26.7
WEA-NY
4 1.272 11.9 1.331 12.4 5.386 40.9 1.268 11.8 1.366 12.9
12 2.448 21.2 2.474 21.6 5.384 40.9 2.510 21.8 2.438 21.5
28 2.674 23.1 2.708 23.4 5.387 40.9 2.797 23.8 2.724 23.1
120 2.713 23.2 2.788 23.5 5.389 40.9 3.063 25.6 2.911 24.6
WEA-SG
4 0.344 1.3 0.346 1.3 0.700 2.5 0.341 1.2 0.346 1.3
12 0.416 1.5 0.421 1.5 0.698 2.5 0.424 1.5 0.425 1.5
28 0.465 1.7 0.473 1.7 0.698 2.5 0.495 1.8 0.471 1.7
120 0.483 1.8 0.496 1.8 0.697 2.5 0.512 1.9 0.485 1.8
Table 3: Performance of Sonnet across the ILI and WEA tasks with different attention modules. Best results are bolded, second best underlined. MAE values are rounded to 3 decimals, sMAPE to 1 for spacing.